# Gravitational lensing by charged black hole with global monopole in strong field limit

### v1

@@ Abstract
We investigate gravitational lensing near a charged black hole with a global monopole in the strong field regime, focusing on two key observables: the angular separation between the first and subsequent images, and the ratio of the flux intensity of the first image to the cumulative flux intensity of all other images. By deriving analytical expressions for these observables, we explore the combined effects of the global monopole and black hole charge both analytically and numerically. Our results reveal that the dependence of the angular separation on charge is intricately tied to the deficit angle caused by the global monopole. In particular, we identify three critical values of the global monopole parameter that determine whether the angular separation increases monotonically, decreases monotonically, or exhibits extrema as the charge varies. A similar complex dependence is found for the flux ratio as a function of the deficit angle. These behaviors contrast sharply with the monotonic changes observed in the absence of either a global monopole or charge. Our findings highlight that the effects of the charge and global monopole on gravitational lensing cannot be described as simple additive contributions. Instead, their combined effects lead to a rich and interdependent behavior that enhances our understanding of strong-field gravitational lensing. While the charge and global monopole are expected to be small in typical astrophysical contexts, the results presented here could be experimentally explored in analogue gravity systems, where these parameters are not constrained. This opens the door to potential experimental verification of the phenomena predicted in this study.
@@ I Introduction
Gravitational lensing is a phenomenon in which the path of light from a distant object is bent by the gravitational field of a massive intervening object (referred to as the lens). This lens could be a black hole, a galaxy, or a galaxy cluster positioned between the source and the observer. The bending of light due to gravity can produce various observable effects, including multiple images of the same object, magnification of the source’s brightness, and the formation of distinctive ring-like structures around the lens, known as Einstein rings. Originally predicted by Albert Einstein within the framework of general relativity Einstein (1936), gravitational lensing has since been extensively observed and confirmed through astronomical observations Walsh et al. (1979).
@@
Gravitational lensing is fundamentally characterized by the deflection angle, which quantifies the bending of light due to gravity. Although the deflection angle itself is not directly observable, it serves as the basis for calculating key observables, such as the angular separation between the first and subsequent images, and the ratio of the flux intensity of the first image to the cumulative flux intensity of all other images. Accurately determining the deflection angle is therefore essential for a comprehensive understanding of gravitational lensing. However, in general situations, directly calculating the deflection angle is challenging. Fortunately, in regimes where the gravitational field is either very weak or extremely strong, approximate methods can be employed. In the weak field limit, where the deflection is small, linearized general relativity provides reliable estimates. Conversely, in strong-field environments, such as near black holes or neutron stars, the gravitational field is strong enough to cause substantial bending of light, sometimes even allowing light to orbit the lens multiple times. In these cases, numerical methods or specialized analytical approximations tailored to highly curved spacetimes are necessary. Here let us note that, gravitational lensing in the strong field limit has been extensively studied in various spacetime backgrounds, including Schwarzschild black holes Bozza (2002), Kerr black holes Bozza (2003); Bozza et al. (2006, 2005), and Reissner-Nordström (RN) black holes Bozza (2002); Tsukamoto (2017), as well as other scenarios Eiroa (2005); Whisker (2005); Tsukamoto et al. (2012); Tsukamoto (2016, 2017); Gyulchev and Yazadjiev (2007); Tsukamoto (2021a, b); Chagoya et al. (2021); Younas et al. (2015); Man and Cheng (2015); Vachher et al. (2024); Bozza (2002); Virbhadra and Ellis (2002). These studies highlight the role of spacetime curvature in determining lensing phenomena.
@@
Apart from spacetime curvature, the topology of spacetime can also influence gravitational lensing. A notable example is a spacetime with a global monopole, a topological defect believed to be formed during phase transitions in the early universe Barriola and Vilenkin (1989). Global monopoles, which arise from phase transitions in systems with self-coupled triplet scalar fields, have garnered significant interest in theoretical physics. Their formation occurs when the original global $O(3)$ symmetry is spontaneously broken to $U(1)$, resulting in a topological defect. These monopoles introduce a deficit solid angle in the surrounding spacetime, leading to distinctive effects, including the bending of light Barriola and Vilenkin (1989) and black hole thermodynamics Yu (1994). For instance, a Schwarzschild black hole that absorbs a global monopole exhibits, in the strong field limit, a reduced relative light flux in its primary image while simultaneously increasing the angular separation between the primary image and the subsequent images Cheng and Man (2011). While previous studies have focused on gravitational lensing effects for black holes with either a global monopole Cheng and Man (2011) or with charge Bozza (2002); Tsukamoto (2017) individually, the combined influence of both charge and a global monopole on gravitational lensing remains unexplored. This raises an intriguing question: Are the effects of a global monopole and charge on black hole gravitational lensing independent, or do they interact in nontrivial ways?
@@
In this paper, we investigate gravitational lensing by a charged black hole with a global monopole in the strong field limit, focusing on two key observables: the angular separation $s$ between the first and subsequent images and the flux intensity ratio $\mathcal{R}$ between the first image and the cumulative flux intensity of all other images. Our study provides a detailed analysis of how these observables depend on the black hole charge and the global monopole parameter, revealing intricate and nontrivial effects. We demonstrate that the angular separation exhibits diverse behaviors as the charge increases, depending on the relationship between the global monopole parameter and three critical values. Specifically, the angular separation may decrease monotonically, increase monotonically, initially decrease and then increase, or even undergo a more complex pattern of initial decrease, subsequent increase, and another decrease at higher charge values. Similarly, a critical charge exists that determines the behavior of the flux ratio with respect to the global monopole parameter. When the charge is below this critical value, the flux ratio increases monotonically with the global monopole parameter, whereas for charge values exceeding the critical value, the flux ratio initially decreases and then increases. These behaviors contrast sharply with the monotonic behaviors in the absence of a global monopole or charge. Our results highlight that the effects of charge and the deficit solid angle due to the presence of the global monopole are not independent but rather interact in a nontrivial manner, significantly altering the gravitational lensing characteristics. While such phenomena require large values of charge and the deficit angle, making them unlikely to be observed in conventional astrophysical black holes, the growing interest in gravitational analogues Barcelo et al. (2005) offers alternative experimental settings. In these analogue systems, effective charge and monopole-like parameters need not be small, making our predicted effects potentially observable in laboratory-based analogue gravity experiments.
@@
The paper is organized as follows. In Sec. II, we present the formula for the deflection angle in the strong field limit and provide general expressions for the observables. In Sec. III, we calculate the deflection angle coefficients in the strong field limit, derive analytical expressions for the observables, and discuss the effects of the charge of black hole and the global monopole. In Sec. IV, we discuss the implications of our findings in analogue gravity systems. Finally, we summarize in Sec. V. Throughout this paper, we adopt natural units $\hbar=c=G_{N}=1$, where $\hbar$ is the reduced Planck constant, $c$ the speed of light, and $G_{N}$ the Newtonian constant of gravitation.
II Deflection angle and observables in the strong field limit
In this section, we first provide a brief review of the calculation of the deflection angle in gravitational lensing within the strong field limit Bozza (2002); Tsukamoto (2017). We outline the key approximations and methods used to derive the deflection angle for light rays passing near the photon sphere of a compact object. Following this, we establish the relationship between the deflection angle and the key observables of gravitational lensing, namely, the angular separation between the first and subsequent images and the flux ratio, by employing the lens equation. This framework allows us to systematically analyze how the presence of charge and a global monopole influences the gravitational lensing characteristics.
Consider an arbitrary static and asymptotically spherically symmetric spacetime with the line element:
$ds^{2}=-A\left(r\right)dt^{2}+B\left(r\right)dr^{2}+C\left(r\right)\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right),$ (2.1)
where $C(r)$ satisfies
$\quad\lim_{r\to\infty}C(r)=r^{2}.$ (2.2)
For simplicity, we assume that both the observer and the light source are located in the equatorial plane of the lens source, i.e. $\theta=\pi/2$. Through this setting, the trajectories of photons are confined to the same plane.
II.1 Physical quantities associated with gravitational lensing
First, we introduce some physical quantities relevant to the study of gravitational lensing.
1. The impact parameter. The impact parameter $b$ is defined as the shortest distance between the straight-line path of a photon if it were unaffected by gravity and the center of the lensing object. The impact parameter $b$ is related to the photon’s angular momentum $L$ and energy $E$ as
$b=\frac{L}{E}.$ (2.3)
@@ II.1 Physical quantities associated with gravitational lensing
2. The turning point. When a photon is emitted from a distant light source, it undergoes gravitational deflection as it passes near a lensing celestial body. There exists a closest distance $r_{0}$ between the photon and the center of the lensing celestial body, and the point $r=r_{0}$ is commonly referred to as the turning point. Based on the conservation of energy and angular momentum, the impact parameter $b$ and the turning point $r_{0}$ are related by the equation
@@
3. The photon sphere. In the vicinity of the event horizon of a black hole, there exists a particularly interesting region known as the photon sphere. On the photon sphere, photons can orbit the black hole along circular trajectories. The radius of the photon sphere, usually denoted as $r_{m}$, can be determined by solving the following equation:
@@
where denotes differentiation with respect to the radial coordinate $r$. This circular orbit is unstable, which means that any small perturbation can cause photons to either fall into the black hole or escape to infinity. Consequently, $r_{m}$ is the closest distance between observable photons and black hole. The photon sphere also defines the critical impact parameter as:
@@
The limit $r_{0}\rightarrow r_{m}$ or $b\rightarrow b_{m}$ is referred to as the strong deflection limit, also known as the strong field limit. Meanwhile, the parameter $b_{m}$ is the smallest impact parameter for observable photons. If the impact parameter of a photon is less than $b_{m}$, it will inevitably fall into the black hole and cannot be observed.
II.2 The deflection angle
Generally, for a photon originating from a source at infinity, undergoing deflection at the turning point $r=r_{0}$, and then escaping towards a distant observer, the deflection angle $\alpha(r_{0})$ can be derived as
@@ II.2 The deflection angle
In the strong field limit, photons undergo multiple orbits near the photon sphere, causing the deflection angle $\alpha$ to diverge. Specifically, $I(r_{0})$ diverges as $r_{0}\rightarrow r_{m}$. To address this issue, a new variable $z$ is introduced, which is defined as
@@
Here, $B(z,r_{0})$, $C(z,r_{0})$ and $R(z,r_{0})$ are obtained by applying the variable substitution in Eq. (2.10) to the original expressions for $B(r)$, $C(r)$ and $R(r)$, respectively.
To isolate the divergent part of the integral (2.11), we partition the integral $I(r_{0})$ into a divergent part $I_{D}(r_{0})$ and a regular part $I_{R}(r_{0})$ as:
@@
$c_{1}(r_{0})=\frac{C_{0}r_{0}}{B_{0}}\left(\frac{C_{0}^{\prime}}{C_{0}}-\frac{A_{0}^{\prime}}{A_{0}}\right)$ (2.16)
@@
$c_{2}(r_{0})=\frac{C_{0}r_{0}}{B_{0}}\left\{\left(\frac{C_{0}^{\prime}}{C_{0}}-\frac{A_{0}^{\prime}}{A_{0}}\right)\left[\left(\frac{C_{0}^{\prime}}{C_{0}}-\frac{A_{0}^{\prime}}{A_{0}}-\frac{B^{{}^{\prime}}_{0}}{B_{0}}\right)r_{0}-3\right]+\frac{r_{0}}{2}\left(\frac{C^{{}^{\prime\prime}}_{0}}{C_{0}}-\frac{A^{{}^{\prime\prime}}_{0}}{A_{0}}\right)\right\}.$ (2.17)
Here, the subscript 00 denotes quantities at $r=r_{0}$. Meanwhile, the regular part $I_{R}(r_{0})$ is given by
@@
Integrating $I_{D}(r_{0})$, we obtain
$I_{D}(r_{0})=\frac{4r_{0}}{\sqrt{c_{2}(r_{0})}}\ln\frac{\sqrt{c_{2}(r_{0})}+\sqrt{c_{1}(r_{0})+c_{2}(r_{0})}}{\sqrt{c_{1}(r_{0})}}.$ (2.19)
In the limit $r_{0}\rightarrow r_{m}$, $c_{1}(r_{0})$ approaches 00 according to Eq. (2.5). Expanding $c_{1}(r_{0})$ around $r_{0}=r_{m}$, we obtain
$c_{1}(r_{0})=\frac{C_{m}r_{m}}{B_{m}}\left(\frac{C_{m}^{{}^{\prime\prime}}}{C_{m}}-\frac{A_{m}^{{}^{\prime\prime}}}{A_{m}}\right)(r_{0}-r_{m})+\mathcal{O}((r_{0}-r_{m})^{2}),$ (2.20)
where the subscript $m$ denotes quantities at $r=r_{m}$. Simultaneously, expanding $b$ around $r_{0}=r_{m}$, we obtain
$b=b_{m}+\frac{1}{4}\sqrt{\frac{C_{m}}{A_{m}}}\left(\frac{C_{m}^{{}^{\prime\prime}}}{C_{m}}-\frac{A_{m}^{{}^{\prime\prime}}}{A_{m}}\right)(r_{0}-r_{m})^{2}+\mathcal{O}((r_{0}-r_{m})^{3}),$ (2.21)
Combining Eqs. (2.19), (2.20) and (2.21), we obtain the divergent part $I_{D}(b)$ in the strong field limit as
$I_{D}(b)=-{\frac{r_{m}}{\sqrt{c_{2}(r_{m})}}}\ln\left(\frac{b}{b_{m}}-1\right)+{\frac{r_{m}}{\sqrt{c_{2}(r_{m})}}}\ln r^{2}_{m}\left(\frac{C_{m}^{{}^{\prime\prime}}}{C_{m}}-\frac{A_{m}^{{}^{\prime\prime}}}{A_{m}}\right)+\mathcal{O}((b-b_{m})\ln(b-b_{m})).$ (2.22)
Finally, by combining Eqs. (2.7), (2.13), (2.18), and (2.22), we can express the deflection angle in the strong field limit as a function of $b$:
@@
respectively.
The above is a brief review of the deflection angle in the strong field limit. The detailed calculation can be found in Refs. Bozza (2002); Tsukamoto (2017). Notably, this method is valid only when a photon sphere exists, i.e., when Eq. (2.5) admits at least one positive solution. Using Eqs. (2.23), (2.24), and (2.25), we can directly calculate the deflection angle of light in the strong field limit. Here, a crucial step is the computation of $I_{R}(r_{m})$. In the case of the Schwarzschild metric, $I_{R}(r_{m})$ admits an exact analytical solution Bozza (2002). Nevertheless, it is noteworthy that not all metrics readily yield such an analytical expression for $I_{R}(r_{m})$.
II.3 Typical observables in gravitational lensing
While the deflection angle of light $\alpha$ and its associated coefficients $\bar{a}$ and $\bar{b}$, are crucial for understanding gravitational lensing, they are not directly measurable. To relate these theoretical parameters to observable quantities, the lens equation serves as a critical tool, converting these quantities into experimentally accessible ones.
Consider the following scenario: A photon emanating from a celestial source ($S$) passes through the gravitational field of a foreground object, termed the lens ($L$). The light ray undergoes deflection, altering its original trajectory and reaching the observer ($O$) at an angle $\theta$, which deviates from the angle $\beta$ that would have been observed in the absence of the lens. The total deflection angle of the light is denoted by $\alpha$. The lens equation can be formulated as follows Bozza et al. (2001):

### v3

# Gravitational lensing by charged black hole with global monopole in the strong field limit

@@ Abstract
We investigate gravitational lensing near a charged black hole with a global monopole in the strong field regime, focusing on the combined effects of the global monopole and black hole charge on key observables in gravitational lensing both analytically and numerically. Our results reveal that the dependence of the angular separation on charge is intricately tied to the deficit angle caused by the global monopole. In particular, we identify three critical values of the global monopole parameter that determine whether the angular separation increases monotonically, decreases monotonically, or exhibits extrema as the charge varies. A similar complex dependence is found for the flux ratio as a function of the deficit angle, and for the magnification of the first relativistic image as a function of charge. These behaviors contrast sharply with the monotonic changes observed in the absence of either a global monopole or charge. Our findings highlight that the effects of the charge and global monopole on gravitational lensing cannot be described as simple additive contributions. Instead, their combined effects lead to a rich and interdependent behavior that enhances our understanding of strong-field gravitational lensing. While the charge and global monopole are expected to be small in typical astrophysical contexts, the results presented here could be experimentally explored in analog gravity systems, where these parameters are not constrained. This opens the door to potential experimental verification of the phenomena predicted in this study.
@@ I Introduction
Gravitational lensing is a phenomenon in which the path of light from a distant object is bent by the gravitational field of a massive intervening object (referred to as the lens). This lens could be a black hole, a galaxy, or a galaxy cluster positioned between the source and the observer. The bending of light due to gravity can produce various observable effects, including multiple images of the same object, magnification of the source’s brightness, and the formation of distinctive ringlike structures around the lens, known as Einstein rings. Originally predicted by Albert Einstein within the framework of general relativity Einstein (1936), gravitational lensing has since been extensively observed and confirmed through astronomical observations Walsh et al. (1979).
@@
Gravitational lensing is fundamentally characterized by the deflection angle, which quantifies the bending of light due to gravity. Although the deflection angle itself is not directly observable, it serves as the basis for calculating key observables, such as the angular separation between the first and subsequent images, the ratio of the flux intensity of the first image to the cumulative flux intensity of all other images, the magnification of the images, and the size of the Einstein rings. Accurately determining the deflection angle is therefore essential for a comprehensive understanding of gravitational lensing. However, in general situations, directly calculating the deflection angle is challenging. Fortunately, in regimes where the gravitational field is either very weak or extremely strong, approximate methods can be employed. In the weak field limit, where the deflection is small, linearized general relativity provides reliable estimates. Conversely, in strong-field environments, such as near black holes or neutron stars, the gravitational field is strong enough to cause substantial bending of light, sometimes even allowing light to orbit the lens multiple times. In these cases, numerical methods or specialized analytical approximations tailored to highly curved spacetimes are necessary. Here let us note that, Virbhadra and Ellis have developed a gravitational lens equation that allows for arbitrarily large deflection angles, which has been applied to analyze gravitational lensing by a Schwarzschild black hole using numerical methods Virbhadra and Ellis (2000). Based on this equation, the position and magnification of relativistic images for a Schwarzschild black hole were derived analytically in Ref. Bozza et al. (2001). Gravitational lensing in the strong field limit for a general spherically symmetric and static spacetime has been investigated by Bozza Bozza (2002) and later by Tsukamoto Tsukamoto (2017). In particular, the Reissner-Nordström (RN) black holes have been studied as one of the specific examples in Refs. Bozza (2002); Tsukamoto (2017). Subsequent studies have explored gravitational lensing in Kerr black holes Bozza (2003); Bozza et al. (2006, 2005), Kerr-Newman black holes Hsieh et al. (2021); Chen et al. (2025), as well as various other curved spacetimes Eiroa (2005); Whisker (2005); Tsukamoto et al. (2012); Tsukamoto (2016); Gyulchev and Yazadjiev (2007); Tsukamoto (2021a, b); Chagoya et al. (2021); Younas et al. (2015); Vachher et al. (2024); Virbhadra and Ellis (2002); Soares et al. (2023). These studies highlight the role of spacetime curvature in determining lensing phenomena.
@@
Apart from spacetime curvature, the topology of spacetime can also influence gravitational lensing Cheng and Man (2011); Man and Cheng (2015); Sharif and Iftikhar (2015); Soares et al. (2024). A notable example is a spacetime with a global monopole, a topological defect believed to be formed during phase transitions in the early universe Barriola and Vilenkin (1989). Global monopoles, which arise from phase transitions in systems with self-coupled triplet scalar fields, have garnered significant interest in theoretical physics. Their formation occurs when the original global $O(3)$ symmetry is spontaneously broken to $U(1)$, resulting in a topological defect. These monopoles introduce a deficit solid angle in the surrounding spacetime, leading to distinctive effects, including the bending of light Barriola and Vilenkin (1989) and black hole thermodynamics Yu (1994). For instance, a Schwarzschild black hole that absorbs a global monopole exhibits, in the strong field limit, a reduced relative light flux in its primary image while simultaneously increasing the angular separation between the primary image and the subsequent images Cheng and Man (2011). While previous studies have focused on gravitational lensing effects for black holes with either a global monopole Cheng and Man (2011) or with charge Bozza (2002); Tsukamoto (2017) individually, the combined influence of both charge and a global monopole on gravitational lensing remains unexplored. This raises an intriguing question: Are the effects of a global monopole and charge on black hole gravitational lensing independent, or do they interact in nontrivial ways?
@@
In this paper, we investigate gravitational lensing by a charged black hole with a global monopole in the strong field limit. In this regime, photons can orbit the lens multiple times, producing a series of relativistic images. We analyze key observables associated with these images, including the angular separation and flux ratio between the first and subsequent images, the angular size of the first Einstein ring, the magnification of the first image, and the time delay between the first and second images. Our analysis reveals intricate and nontrivial dependencies of these observables on the black hole charge and the global monopole parameter. In particular, the angular separation exhibits diverse behaviors as the charge increases, depending on the relationship between the global monopole parameter and three critical values. Specifically, the angular separation may decrease monotonically, increase monotonically, initially decrease and then increase, or even undergo a more complex pattern of initial decrease, subsequent increase, and another decrease at higher charge values. Similarly, a critical charge exists that determines the behavior of the flux ratio with respect to the global monopole parameter. When the charge is below this critical value, the flux ratio increases monotonically with the global monopole parameter, whereas for charge values exceeding the critical value, the flux ratio initially decreases and then increases. Additionally, we identify two critical values of the global monopole parameter that determine whether the magnification of the first relativistic image monotonically decreases, increases, or exhibits nonmonotonic dependence on the charge. These rich behaviors stand in stark contrast to the simpler, monotonic trends observed in the absence of either charge or a global monopole. Our results highlight that the effects of charge and the deficit solid angle due to the presence of the global monopole are not independent but rather interact in a nontrivial manner, significantly altering the gravitational lensing characteristics. Furthermore, we numerically calculate the observables for scenarios in which the central black holes in the Milky Way, M87, and M104 galaxies act as gravitational lenses, comparing configurations with and without charge and a global monopole. While the resulting differences are small and likely challenging to detect in astrophysical contexts, the growing interest in gravitational analogs Barcelo et al. (2005) offers alternative experimental settings. In such analog systems, effective charge and monopolelike parameters need not be small, making our predicted effects potentially observable in laboratory-based analog gravity experiments.
@@
The paper is organized as follows. In Sec. II, we derive the deflection angle for an RN black hole with a global monopole in the strong field limit. In Sec. III, we analyze the effects of charge and global monopole on typical gravitational lensing observables both analytically and numerically. In Sec. IV, we discuss the potential detectability of these effects in astrophysical settings and their implications in analog gravity systems. Finally, we summarize in Sec. V. Throughout this paper, we adopt natural units $\hbar=c=G_{N}=1$, where $\hbar$ is the reduced Planck constant, $c$ is the speed of light, and $G_{N}$ is the Newtonian constant of gravitation.
II Deflection angle in strong field limit
We now investigate the gravitational lensing of an RN black hole with a global monopole. Its line element is given by Barriola and Vilenkin (1989); Guendelman and Rabinowitz (1991)
$ds^{2}=-A(r)dt^{2}+B(r)dr^{2}+C(r)\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right),$ (2.1)
where
$A(r)=B^{-1}(r)=1-8\pi\gamma^{2}-\frac{2M}{r}+\frac{Q^{2}}{r^{2}},\quad C(r)=r^{2}.$ (2.2)
$\epsilon=1-8\pi\gamma^{2}=1-\frac{\Delta}{4\pi},$ (2.3)
which ranges from $\epsilon=0$ (corresponding to a maximum deficit solid angle $\Delta=4\pi$) to $\epsilon=1$ (when no monopole is present, i.e., $\Delta=0$), and the charge parameter $\eta$ as
$\eta=Q^{2}/M^{2}.$ (2.4)
This parameter quantifies the effect of charge, with values ranging from $\eta=0$ (uncharged black hole) to $\eta=1$ (extremal black hole in standard RN spacetime). To ensure that the spacetime does not develop a naked singularity, we impose the constraint $|Q|\leq M/\sqrt{\epsilon}$. For generality, and to avoid naked singularities for any $\epsilon$, we restrict the charge to $|Q|\leq M$, leading to the constraint $\eta\in[0,1]$. With these dimensionless parameters, the functions $A(r)$ and $B(r)$ in the metric can be rewritten as
$A(r)=B^{-1}(r)=\epsilon-\frac{2M}{r}+\frac{\eta M^{2}}{r^{2}}.$ (2.5)
In the following, we calculate the deflection angle in the strong field limit. We start by introducing the relevant physical quantities associated with gravitational lensing and provide a precise definition of the strong field regime.
@@ II Deflection angle in strong field limit
When a photon is emitted from a distant light source, it undergoes gravitational deflection as it passes near a lensing celestial body. The impact parameter $b$ is defined as the shortest distance between the straight-line path of a photon if it were unaffected by gravity and the center of the lensing object. In addition, there exists a closest distance $r_{0}$ between the photon and the center of the lensing celestial body, and the point $r=r_{0}$ is commonly referred to as the turning point. Based on the conservation of energy and angular momentum, the impact parameter $b$ and the turning point $r_{0}$ are related by the equation
@@
In the vicinity of the event horizon of a black hole, there exists a particularly interesting region known as the photon sphere. On the photon sphere, photons can orbit the black hole along circular trajectories. The radius of the photon sphere, usually denoted as $r_{m}$, can be determined by solving the following equation Claudel et al. (2001):
@@
where denotes differentiation with respect to the radial coordinate $r$. This circular orbit is unstable, which means that any small perturbation can cause photons to either fall into the black hole or escape to infinity. Consequently, $r_{m}$ is the closest distance between observable photons and black hole. The photon sphere also defines the critical impact parameter as Bozza (2002); Tsukamoto (2017)
@@
For an RN black hole with a global monopole, the radius of the photon sphere and the critical impact parameter are given by
$r_{m}\left(\epsilon,\eta\right)=\frac{3+\sqrt{9-8\epsilon\eta}}{2\epsilon}M,$ (2.9)
and
$b_{m}\left(\epsilon,\eta\right)=\frac{\bar{r}_{m}^{2}}{\sqrt{\epsilon\bar{r}_{m}^{2}-2\bar{r}_{m}+\eta}}M,$ (2.10)
respectively, where $\bar{r}_{m}\left(\epsilon,\eta\right)=r_{m}\left(\epsilon,\eta\right)/M=(3+\sqrt{9-8\epsilon\eta})/2\epsilon$ is the dimensionless radius of the photon sphere. The limit $r_{0}\rightarrow r_{m}$ or $b\rightarrow b_{m}$ is referred to as the strong deflection limit, also known as the strong field limit. Meanwhile, the parameter $b_{m}$ is the smallest impact parameter for observable photons. If the impact parameter of a photon is less than $b_{m}$, it will inevitably fall into the black hole and cannot be observed.
Generally, for a photon originating from a source at infinity, undergoing deflection at the turning point $r=r_{0}$, and then escaping toward a distant observer, the deflection angle $\alpha(r_{0})$ can be derived as
@@
In the strong field limit, photons undergo multiple orbits near the photon sphere, causing the deflection angle $\alpha$ to diverge. Specifically, $I(r_{0})$ diverges as $r_{0}\rightarrow r_{m}$. To address this issue, a new variable $z$ is introduced, which is defined as Tsukamoto (2017)
@@
Here, $B(z,r_{0})$, $C(z,r_{0})$, and $R(z,r_{0})$ are obtained by applying the variable substitution in Eq. (2.14) to the original expressions for $B(r)$, $C(r)$, and $R(r)$, respectively.
To isolate the divergent part of the integral (2.15), we partition the integral $I(r_{0})$ into a divergent part $I_{D}(r_{0})$ and a regular part $I_{R}(r_{0})$ as
@@
$c_{1}(r_{0})=\frac{C(r_{0})r_{0}}{B(r_{0})}\left(\frac{C^{\prime}(r_{0})}{C(r_{0})}-\frac{A^{\prime}(r_{0})}{A(r_{0})}\right)$ (2.20)
@@
$c_{2}(r_{0})=\frac{C(r_{0})r_{0}}{B(r_{0})}\left\{\left(\frac{C^{\prime}(r_{0})}{C(r_{0})}-\frac{A^{\prime}(r_{0})}{A(r_{0})}\right)\left[\left(\frac{C^{\prime}(r_{0})}{C(r_{0})}-\frac{A^{\prime}(r_{0})}{A(r_{0})}-\frac{B^{\prime}(r_{0})}{B(r_{0})}\right)r_{0}-3\right]+\frac{r_{0}}{2}\left(\frac{C^{{}^{\prime\prime}}(r_{0})}{C(r_{0})}-\frac{A^{{}^{\prime\prime}}(r_{0})}{A(r_{0})}\right)\right\}.$ (2.21)
Meanwhile, the regular part $I_{R}(r_{0})$ is given by
@@
After some calculations, the deflection angle $\alpha$ in the strong field limit is derived as Tsukamoto (2017)
@@
respectively, with the subscript $m$ denoting quantities evaluated at at $r=r_{m}$. For a detailed derivation, see Ref. Tsukamoto (2017).
Substituting the metric functions for an RN black hole with a global monopole Eq. (2.5) into Eqs. (2.24) and (2.25), we obtain the coefficients $\bar{a}$ and $\bar{b}$ which determine the deflection angle in the strong field limit as
$\bar{a}(\epsilon,\eta)=\frac{\bar{r}_{m}\left(\epsilon,\eta\right)}{\sqrt{\epsilon\bar{r}_{m}^{2}\left(\epsilon,\eta\right)-2\eta}},$ (2.26)
and
$\bar{b}(\epsilon,\eta)=\bar{a}(\epsilon,\eta)\ln\left[\frac{8\left(3\bar{r}_{m}(\epsilon,\eta)-4\eta\right)^{3}}{\bar{r}_{m}^{2}(\epsilon,\eta)\left(\bar{r}_{m}(\epsilon,\eta)-\eta\right)^{2}}\left(2\sqrt{\bar{r}_{m}(\epsilon,\eta)-\eta}-\sqrt{3\bar{r}_{m}(\epsilon,\eta)-4\eta}\right)^{2}\right]-\pi,$ (2.27)
respectively. When the global monopole is absent ($\epsilon=1$), the coefficients $\bar{a}$ and $\bar{b}$ recover the results for the RN black hole, as derived in Ref. Tsukamoto (2017). Similarly, in the absence of charge ($\eta=0$), the coefficients $\bar{a}$ and $\bar{b}$ reduce to the expressions obtained for a Schwarzschild black hole with a global monopole, as given in Ref. Cheng and Man (2011). Furthermore, in the absence of both charge and the global monopole ($\eta=0$, $\epsilon=1$), the coefficients $\bar{a}$ and $\bar{b}$ match the results from Refs. Bozza (2002); Tsukamoto (2017) for a Schwarzschild black hole. These consistency checks further validate our approach, ensuring that the inclusion of both charge and a global monopole provides a unified framework that naturally extends existing results.
III Interdependence of gravitational lensing observables on charge and global monopole
While the deflection angle of light $\alpha$ and its associated coefficients $\bar{a}$ and $\bar{b}$ are crucial for understanding gravitational lensing, they are not directly measurable. To relate these theoretical parameters to observable quantities, the lens equation serves as a critical tool, converting these quantities into experimentally accessible ones.
Consider the following scenario: A photon emanating from a celestial source ($S$) passes through the gravitational field of a foreground object, termed the lens ($L$). The light ray undergoes deflection, altering its original trajectory and reaching the observer ($O$) at an angle $\theta$, which deviates from the angle $\beta$ that would have been observed in the absence of the lens. The total deflection angle of the light is denoted by $\alpha$. The lens equation can be formulated as follows Virbhadra and Ellis (2000):